OpenStudy (anonymous):

If a line crosses the y-axis at (0, 1) and has a slope of 45, what is the equation of the line. Ok so it specifies that you need to find the equation for the slope and the y intercept. Is it either 5y+4x or 5y-4x=5 i say it's positive due to the slope goes uphill. What do you think ?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

going*

OpenStudy (whpalmer4):

The slope is 45? 45 degrees, or an actual slope m = 45?

OpenStudy (anonymous):

sorry that was suppose to be a fraction lol 4/5

OpenStudy (whpalmer4):

ah, yeah, that makes a difference :-) \[\text{slope }m = \frac{4}5\]\[(x_0,y_0) = (0,1)\] \[y-y_0 = m(x-x_0)\]\[y - 1 = \frac{4}{5}(x-0)\]That's the point-slope equation, rearrange to the form you want. Alternatively, \[m = \frac{4}{5}\]\[b=1\]\[y = mx+b = \frac{4}{5}x+1\]That is the slope-intercept equation, rearrange to the form you want. The first one is a bit more general, as you can use it with any point, not just the y-intercept.

OpenStudy (anonymous):

So i can rearrange like so ?5y + 4x = 5

OpenStudy (whpalmer4):

if you want standard form, I'd multiply by 5: \[5y = 5*\frac{4}{5}x+5*1\]\[5y = 4x + 5\]\[5y-4x=5\]

OpenStudy (whpalmer4):

remember, if you drag a term across the equals sign like we did with \(4x\), you need to change its sign.

OpenStudy (anonymous):

ThThank you for that explaintion, i knew when to switch signs